# Homological shifts of a complementary edge ideal

### v1

@@ Abstract.
The homological shift algebra and the projective dimension function of complementary edge ideals are investigated. Let $G$ be a connected graph, and let $I$ be its complementary edge ideal. For bipartite graphs $G$, we show that the projective dimension of $I^{s}$ increases strictly with $s$ until reaching its maximum value. For trees and cycles, explicit expressions for the projective dimension of $I^{s}$ are provided, along with detailed descriptions of their homological shift algebras. In particular, it is shown that the $i$-th homological shift algebra of such ideals is generated in degree at most $i$. Additionally, we prove that if $G$ is a tree, then the homological shift ideal $\mathrm{HS}_{i}(I^{i})$, when divided by a suitable monomial, becomes a Veronese-type ideal, and every Veronese-type ideal arises in this manner.
2020 Mathematics Subject Classification. Primary 13A02, 13C15; Secondary 05E40, 05C05, 05C38
@@
Keywords: projective dimension, tree, cycle, homological shift ideal, complementary edge ideal
Introduction
Let $S={\mathbb{K}}[x_{1},\ldots,x_{n}]$ be the polynomial ring in $n$ variables over a field $\mathbb{K}$}, and let $I\subset S$ be a non-zero, squarefree monomial ideal generated in degree $d$. Fröberg showed in [20] that if $d=2$, then $I$ has a linear resolution precisely when $I$ is the edge ideal of a chordal graph. Subsequently, Herzog, Hibi and Zheng proved in [9] that $I$ has a linear resolution if and only if all powers $I^{k}$ have linear quotients (also see [22, Theorem 3.2]). A monomial ideal $I$ is fully supported if for every $1\leq k\leq n$, the variable $x_{k}$ divides some minimal monomial generator of $I$. More recently, Ficarra and Moradi demonstrated in [15, Corollary 3.2] that if $I$ is fully supported in $S$ and is generated in degree d=n2d=n-2, then $I$ has a linear resolution if and only if $I^{k}$ has linear quotients for all $k\geq 1$. For any non-zero squarefree monomial ideal in degree n2n-2, it can be regarded as the complementary edge ideal of a finite simple graph $G$ with $n$ vertices. Let $G$ be a simple graph with with $n$ vertices, and let $E(G)$ be its edge set. We then define the complementary edge ideal of $G$ as
$I_{c}(G)=\left(\frac{{\mathbf{x}}_{[n]}}{x_{i}x_{j}}:\ \{i,j\}\in E(G)\right),$
where ${\mathbf{x}}_{[n]}=\prod\limits_{i=1}^{n}x_{i}$ and $n]=\{1,\ldots,n\}$. Furthermore, Ficarra and Moradi also showed in [15, Corollary 3.2] that $I_{c}(G)$ has a linear resolution if and only if $G$ is connected.
The complementary edge ideal of a graph was independently studied in 2025 in [16] and [28]. Hibi et al., as well as Ficarra and Moradi independently characterized sequentially Cohen-Macaulay, Cohen-Macaulay and Gorenstein complementary edge ideals. Furthermore, Ficarra and Moradi also characterized nearly Gorenstein and matroidal complementary edge ideals. They also computed the regularity and the asymptotic depth of powers of complementary edge ideals, as well as the index of their depth stability.
@@ Introduction
To study the syzygies and resolutions of monomial ideals, we use the homological shift ideal. This concept first appeared implicitly in Miller and Sturmfels’s 2005 book [30, Theorem 2.18] and was formally defined by Herzog et al. in 2020 [25].
@@
Let ${\mathbf{a}}=(a_{1},\ldots,a_{n})\in{\mathbb{Z}}_{\geq 0}^{n}$ and ${\mathbf{x}}^{{\mathbf{a}}}=x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$. Given a monomial ideal $I\subset S$, the $i$-th homological shift ideals of $I$ is defined as for all $i\geq 0$:
$\operatorname{HS}_{i}(I)=({\mathbf{x}}^{{\mathbf{a}}}\mid\operatorname{Tor}_{i}^{S}({\mathbb{K}},I)_{{\mathbf{a}}}\neq 0).$
Note that $\mbox{HS}_{0}(I)=I$. Homological shift ideals encode critical multigraded data arising from the syzygies of $I$, see [2, 3, 10, 11, 12, 13, 14, 17, 18, 19, 25, 26, 31, 33]. A major motivation for studying homological shift ideals stems from the Bandari-Bayati-Herzog conjecture that being polymatroidal is a homological shift property, which has been answered positively in [6] recently. Inspired by conjectures from [2] and [25], Ficarra and Qureshi introduced the $i$-th homological shift algebra $\mathrm{HS}_{i}(\mathcal{R}(I))$ in [19]. By definition, this algebra is given by
@@
When $I$ has linear powers, they showed that $I\cdot\mathrm{HS}_{i}(I^{s})\subseteq\mathrm{HS}_{i}(I^{s+1})$ for all $s\geq 0$, which implies that $\mathrm{HS}_{i}(\mathcal{R}(I))$ is a graded $\mathcal{R}(I)$-module. Furthermore, it is finitely generated (see [19, Theorem 1.4]). The generating degree of homological shift algebras for a polymatroidal ideal has been investigated in [13]. Ficarra and Lu conjecture that, if $I$ is a polymatroidal ideal, then $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degrees $\leq i$ as an $\mathcal{R}(I)$-module for all $i>0$ (see [13, Conjecture 5.1]). They proved that this conjecture is true for special values of $i$ such as $i=1$, as well as for special polymatroidal ideals, including principal Borel ideals and polymatroidal ideals that satisfy the strong exchange property. Matroidal edge ideals are another example (see [5, 13]).
@@
In this paper, we focus on the algebraic and homological properties of the complementary edge ideal of a graph and its powers. Specifically, we focus on the homological shift ideal and the projective dimension of powers of the complementary edge ideals of trees and cycles. Let $\mathrm{pd}(-)$ denote the projective dimension of an ideal -. Our results are as follows:
@@
Let $G$ be a simple graph and let $I_{c}(G)$ be its complementary edge ideals.
@@
    If $G$ is an even cycle with $2n$ vertices, then
    $\mathrm{pd}(I_{c}(G)^{s})=\left\{\begin{array}[]{ll}2s,&\hbox{\text{ if\ } $s\in[n-1]$;}\\ 2n-2,&\hbox{\text{ if\ } $s\geq n-1$.}\end{array}\right$.
@@
    If $G$ is an odd cycle with $2n+1$ vertices, then
    $\mathrm{pd}(I_{c}(G)^{s})=\left\{\begin{array}[]{ll}2s,&\hbox{\text{ if\ } $s\in[n]$;}\\ 2n,&\hbox{\text{ if\ } $s>n$.}\end{array}\right$.
@@
Let $G$ be a connected bipartite graph on vertex set $[n]$. Then $\mathrm{pd}(I_{c}(G)^{s})$ strictly increases until it reaches the maximal value n2n-2.
Let $\alpha$ denote the monomial $\prod\limits_{i=1}^{n}x_{i}$.
@@
Let $G$ be a connected graph and let $I=I_{c}(G)$ be its complementary edge ideal. Then for all $k\geq 0$, we have
$\mathrm{HS}_{1}(I^{k+1})=I^{k}\cdot\mathrm{HS}_{1}(I).$
@@
Let $G$ be a tree with $n$ vertices, labelled as in Remark 3.5, and let $I=I_{c}(G)$ be its complementary edge ideal. Define a map $\phi:[n-1]\rightarrow[n]$ such that $\phi(i)$ is the unique $j\in[n]$, with $j>i$ for which $x_{i}x_{j}\in E(G)$. Then, for any $1\leq i\leq s$, we have
$\mathrm{HS}_{i}(I^{s})=I^{s-i}\cdot\left(\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}:\ F\subseteq[n-2]\text{\ and\ }|F|=i\right).$
As an immediate consequence of this result, if $G$ is tree, then the $i$-th homological shift algebra $\mathrm{HS}_{i}(\mathcal{R}(I_{c}(G)))$ is generated in degree $i$ for all $1\leq i\leq n-2$.
@@
Let $G$ be a tree let $I=I_{c}(G)$ be its complementary edge ideal. Then $\operatorname{HS}_{i}(I_{c}(G)^{i})$, when divided by a suitable monomial, is an ideal of Veronese type. Furthermore. every ideal of Veronese type arises in this manner.
Studying the homological shift algebra of the complementary edge ideal of a cycle is much more complicated. Regarding this topic, our results are as follows:
@@
$\mathrm{HS}_{i}(I^{s})=\sum_{\begin{subarray}{c}\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor\end{subarray}}\left(\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i-2k\end{subarray}}\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}\right)$
@@
Let $G$ be a cycle and let $I_{c}(G)$ be its complementary edge ideal, then $\mathrm{HS}_{i}(\mathcal{R}(I_{c}(G)))$ is generated in degree at most iqi-q for all $i>0$. Here, $q=\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}$.
The article is organized as follows: Section 1 provides the definitions and basic facts that will be used throughout the paper. In Section 2, we present a theorem on the homological shift ideal of independent significance. In Section 3, we compute the projective dimension of powers of the complementary edge ideals of trees and cycles. Furthermore, if $G$ is a connected bipartite graph, we prove that $\mathrm{pd}(I_{c}(G)^{s})$ strictly increases until it reaches its maximum value. In Section 4, we present the explicit expressions of homological shift ideals of powers for the complementary edge ideals of trees and cycles.
1. Preliminary
This section provides the definitions and basic facts used throughout the paper.
For a vector ${\mathbf{a}}=(a_{1},\ldots,a_{n})\in{\mathbb{Z}}^{n}$, we set ${\mathbf{x}}^{{\mathbf{a}}}=x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$. For a subset $F\subseteq[n]$, we set ${\mathbf{x}}_{F}=\prod_{i\in F}x_{i}$. Given positive integers $k,n$ with $k\leq n$, let $[k,n]$ denote the set $\{k,k+1,\ldots,n\}$. In particular, if $k=1$, $[k,n]$ is denoted by $[n]$.

### v5

# Homological shifts of powers of a complementary edge ideal

@@ Abstract.
The homological shift algebra and the projective dimension function of a complementary edge ideal are investigated. Let $G$ be a connected graph, and let $I$ be its complementary edge ideal. For bipartite graphs $G$, we show that the projective dimension of $I^{s}$ increases strictly with $s$ until reaching its maximum value. For trees and cycles, explicit expressions for the projective dimension of $I^{s}$ are provided, along with precise formulas of their homological shift ideals. In particular, it is shown that the $i$-th homological shift algebra of such ideals is generated in degree at most $i$. We prove that if $G$ is a tree, then the homological shift ideal $\mathrm{HS}_{i}(I^{i})$, when divided by a suitable monomial, is a Veronese-type ideal. Moreover, every Veronese-type ideal can be realized in this manner.
2020 Mathematics Subject Classification. Primary 05E40; Secondary 13A02, 13C15
@@
Keywords: projective dimension, tree, homological shift algebra, complementary edge ideal
1. Introduction
@@
Let $S:=\mathbb{K}[x_{1},\ldots,x_{n}]$ denote a polynomial ring over a field $\mathbb{K}$. In the study of syzygies and free resolutions of monomial ideals in $S$, the concept of homological shift ideals serves as a powerful algebraic tool. While this concept first appeared implicitly in the 2005 monograph of Miller and Sturmfels [30], it was not formally defined until Herzog et al. in 2021 [23].
@@
Given a vector ${\mathbf{a}}=(a_{1},\ldots,a_{n})\in{\mathbb{Z}}_{\geq 0}^{n}$, let ${\mathbf{x}}^{{\mathbf{a}}}$ denote the monomial $x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$ in $S$. For a monomial ideal $I\subset S$, we adopt the definition introduced in [23] and define the $i$-th homological shift ideal of $I$ (for each integer $i\geq 0$) as
$\operatorname{HS}_{i}(I):=\left({\mathbf{x}}^{{\mathbf{a}}}\mid\operatorname{Tor}_{i}^{S}(\mathbb{K},I)_{{\mathbf{a}}}\neq 0\right).$
A central question in the study of homological shift ideals concerns the identification of which properties of a monomial ideal $I$ are preserved by all its homological shift ideals $\mathrm{HS}_{i}(I)$. A key property to investigate in this regard is polymatroidality: if $I$ is a polymatroidal ideal, do all $\mathrm{HS}_{i}(I)$ inherit this property? Partial results addressing this question have appeared in the literature: Bayati [2] proved that all $\mathrm{HS}_{i}(I)$ of a matroidal ideal $I$ remain matroidal; Herzog et al. [23] showed that the homological shift ideals of Veronese-type ideals are polymatroidal; Ficarra [7] established that $\mathrm{HS}_{1}(I)$ is polymatroidal whenever $I$ is polymatroidal. Recently, a complete and positive answer to this question has been established in [6]. Another property of interest pertains to linear quotients: if $I$ admits linear quotients, do all $\mathrm{HS}_{i}(I)$ inherit this property? This holds for $i=1$, as demonstrated in [11]; however, it fails to hold in general for $i\geq 2$. Multiple counterexamples have been constructed in the literature to confirm this fact, and [29, Proposition 3.5] provides one such instance.
Of equal, if not greater, significance to the above problem is the explicit construction of homological shift ideals for specific families of monomial ideals. For instance, if $I$ is an equigenerated square-free Borel ideal, then $\mathrm{HS}_{1}(I)$ has been explicitly characterized in [25]; if $I$ is the edge ideal of a co-chordal graph or a Veronese-type ideal, the structure of $\mathrm{HS}_{i}(I)$ has been rigorously determined in [23]; if $I$ is the generalized co-letterplace ideal of a poset, $\mathrm{HS}_{i}(I)$ has been explicitly described in [29]. Additional advances in the study of homological shift ideals can be found in [3, 9, 31, 32].
To study the asymptotic properties of homological shift ideals, Ficarra and Qureshi introduced the $i$-th homological shift algebra $\mathrm{HS}_{i}(\mathcal{R}(I))$ in [16]. By definition, this abelian group structure is given, for all $i\geq 1$, by:
@@
When $I$ has linear powers (meaning $I^{s}$ admits a linear resolution for all $s\geq 1$), Ficarra and Qureshi showed in [16, Proposition 1.2] that $I\cdot\mathrm{HS}_{i}(I^{s})\subseteq\mathrm{HS}_{i}(I^{s+1})$ for all $s\geq 0$. This containment implies $\mathrm{HS}_{i}(\mathcal{R}(I))$ is a graded $\mathcal{R}(I)$-module. Furthermore, by [16, Theorem 1.4], this graded module is finitely generated. The generating degree of homological shift algebras over $\mathcal{R}(I)$ has been investigated in [15] and [10]. In [15], it was shown that if $I$ is the edge ideal of a simple graph, then an explicit expression for $\mathrm{HS}_{1}(I^{s})$ can be derived, from which it follows that $\mathrm{HS}_{1}(\mathcal{R}(I))$ is generated in degree $1$. In [10], Ficarra and Lu proposed the following conjecture: if $I$ is a polymatroidal ideal, then $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degrees at most $i$ as an $\mathcal{R}(I)$-module for all $i>0$ (see [10, Conjecture 5.1]). They verified this conjecture for special cases, including $i=1$ and specific classes of polymatroidal ideals such as principal Borel ideals, polymatroidal ideals satisfying the strong exchange property, and matroidal edge ideals. Another class of polymatroidal ideals satisfying the conjecture was provided in [5]. It is natural to reformulate [10, Conjecture 5.1] in a more general sense:
Conjecture 1.1.
If $I$ is a monomial ideal for which $I^{s}$ admits linear quotients for all $s\geq 1$, then $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degrees at most $i$ over $\mathcal{R}(I)$ for all $i\geq 1$.
Let $I\subset S$ be a non-zero square-free monomial ideal generated in degree $d$. Note that $1\leq d\leq n$. If $d=1$ or $d=n$, II has a simple structure, leaving no non-trivial problems for further investigation.
Within the range $2\leq d\leq n-2$, two distinguished values of $d$ ($d=2$ and d=n2d=n-2) exhibit distinctive behavior relative to other degrees, particularly regarding the linear resolutions and linear quotient property of powers of such ideals.
For $d=2$, Fröberg’s classic result [17] asserts that $I$ has a linear resolution if and only if $I$ is the edge ideal of a co-chordal graph. Herzog, Hibi, and Zheng further extended this equivalence to powers of such ideals [20, Theorem 3.2]: for square-free monomial ideals of degree 2, $I$ has a linear resolution if and only if $I^{s}$ has linear quotients for all $s\geq 1$.
A monomial ideal $I\subset S$ is fully supported if every variable $x_{k}$ ($1\leq k\leq n$) divides some minimal monomial generator of $I$. For the fully supported square-free monomial ideals of degree d=n2d=n-2, Ficarra and Moradi [12, Corollary 3.2] and Hibi et al. [27, Theorem 2.2] independently established a parallel characterization: $I$ has a linear resolution if and only if $I^{s}$ has linear quotients for all $s\geq 1$.
By contrast, for intermediate degrees $3\leq d\leq n-3$, [12, Theorem D] shows that there exists a fully supported square-free monomial ideal $I\subset S$ of degree $d$ such that $I$ has linear quotients but $I^{2}$ fails to admit a linear resolution. Moreover, [12, Theorem C] shows that there exists a fully supported square-free monomial ideal $I\subset S$ generated in degree $d$ that fails to admit a linear resolution when $\operatorname{char}(\mathbb{K})=2$, yet it admits one otherwise.
For $d=2$, every square-free monomial ideal of degree $d$ is an edge ideal, a class with a long history of extensive investigation. In sharp contrast, the class of fully supported square-free monomial ideals of degree d=n2d=n-2 has only recently garnered considerable attention in combinatorial commutative algebra, owing to independent work by two teams in 2025: Ficarra and Moradi [12, 13, 14] and Hibi et al. [27]. Notably, every such ideal is the complementary edge ideal of a finite simple graph $G$ on $n$ vertices, defined as follows. Let $G$ be a simple graph with vertex set $n]=\{1,2,\dots,n\}$ and edge set $E(G)$. Set ${\mathbf{x}}_{[n]}:=\prod_{i=1}^{n}x_{i}$; the complementary edge ideal of $G$ is then
$I_{c}(G):=\left(\frac{{\mathbf{x}}_{[n]}}{x_{i}x_{j}}:\{i,j\}\in E(G)\right).$
Ficarra and Moradi [12, Corollary 3.2] and Hibi et al. [27, Theorem 2.2] further proved that $I_{c}(G)$ has a linear resolution if and only if $G$ is connected.
Beyond characterizing linear resolution conditions for these ideals and their powers, the two teams independently established the sequentially Cohen-Macaulay, Cohen-Macaulay, and Gorenstein properties of $I_{c}(G)$. Ficarra and Moradi extended these findings by characterizing the nearly Gorenstein and matroidal subclasses of complementary edge ideals. Additionally, they computed key homological invariants of the powers of these ideals, including regularity and depth, among others.
@@
In this paper, we investigate the algebraic and homological properties of the complementary edge ideal of a simple graph and its powers. Specifically, we focus on determining the homological shift algebra and the projective dimension function of the complementary edge ideals of trees and cycles, with particular emphasis on their links to Conjecture 1.1.
Our first result is not directly related to the other results in this paper, but is of independent interest.
@@
(Theorem 3.1) Let $I\subseteq S=\mathbb{K}[x_{1},\ldots,x_{n}]$ be a monomial ideal with a $d$-linear resolution. Then for all $i\in[n]$, we have
$\mathrm{HS}_{i}(I)+\mathrm{HS}_{i-1}(\mathfrak{m}I)=\mathfrak{m}^{[i]}I.$
Next, we investigate the projective dimension of powers of a complementary edge ideal. Let $\mathrm{pd}(-)$ denote the projective dimension of an ideal.
Theorem 1.3.
(Theorem 4.4 and Theorem 4.8) Let $G$ be a simple graph and let $I_{c}(G)$ be its complementary edge ideal.
@@
    If $G$ is an even cycle with $2m$ vertices, then
    $\mathrm{pd}(I_{c}(G)^{s})=\left\{\begin{array}[]{ll}2s,&\hbox{\text{ if\ } $s\in[m-1]$;}\\ 2m-2,&\hbox{\text{ if\ } $s\geq m-1$.}\end{array}\right$.
@@
    If $G$ is an odd cycle with $2m+1$ vertices, then
    $\mathrm{pd}(I_{c}(G)^{s})=\left\{\begin{array}[]{ll}2s,&\hbox{\text{ if\ } $s\in[m]$;}\\ 2m,&\hbox{\text{ if\ } $s>m$.}\end{array}\right$.
@@
(Proposition 4.10) Let $G$ be a connected graph on vertex set $[n]$. Then $\mathrm{pd}(I_{c}(G)^{s})$ is strictly increasing in $s$ before reaching n2n-2.
If $G$ is bipartite, $\mathrm{pd}(I_{c}(G)^{s})$ remains constant at n2n-2 for all $s$ beyond the point where it first reaches n2n-2. If $G$ is non-bipartite, the maximal value of $\mathrm{pd}(I_{c}(G)^{s})$ over $s\geq 1$ is n1n-1. It remains undetermined whether $\mathrm{pd}(I_{c}(G)^{s})$ jumps to n1n-1 right after attaining n2n-2.
Given an ideal $I$, we let $\mathrm{dstab}(I)$ denote the smallest integer $k$ such that $\mathrm{pd}(I^{s})$ remains constant for all integers $s\geq k$.
@@
(Corollary 4.11 and Theorem 4.13) Let $G$ be a connected graph with $n$ vertices. Then
$\mathrm{dstab}(I_{c}(G))\leq n-2.$
The main theme of this paper is to describe the homological shift algebra of a complementary edge ideal.
@@
(Theorem 5.1) Let $G$ be a connected graph on vertex set $[n]$ and let $I=I_{c}(G)$ be its complementary edge ideal. Then the 1-st homological shift algebra $\operatorname{HS}_{1}(\mathcal{R}(I))$ is generated in degree $1$ as a graded $\mathcal{R}(I)$-module.
Let $\alpha$ denote the square-free monomial $x_{1}x_{2}\cdots x_{n}$.
@@
(Theorem 5.2) Let $G$ be a tree with $n$ vertices, labelled as in Remark 4.5, and let $I=I_{c}(G)$ be its complementary edge ideal. Define a map $\phi:[n-1]\rightarrow[n]$ such that $\phi(j)$ is the unique $k\in[n]$, with $k>j$ for which $x_{j}x_{k}\in E(G)$. Then, for any $1\leq i\leq s$, we have
$\mathrm{HS}_{i}(I^{s})=I^{s-i}\cdot\left(\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}\middle|\;F\subseteq[n-2]\text{\ and\ }|F|=i\right).$
In particular, the $i$-th homological shift algebra $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated exactly in degree $i$ for all $1\leq i\leq n-2$.
Let $d$ be a positive integer, and $\boldsymbol{c}=(c_{1},\dots,c_{n})\in\mathbb{Z}_{\geq 0}^{n}$ be a given non-negative integer vector. The Veronese-type ideal of $S$ associated with $d$ and $\boldsymbol{c}$ is a monomial ideal generated by all monomials in $S$ of total degree $d$ whose exponents are componentwise bounded by $\boldsymbol{c}$. In symbolic form:
$I_{\boldsymbol{c},d}=\Bigl(x_{1}^{a_{1}}x_{2}^{a_{2}}\cdots x_{n}^{a_{n}}\ \Bigm|\ \sum_{i=1}^{n}a_{i}=d,\quad 0\leq a_{i}\leq c_{i}\ \text{for all}\ 1\leq i\leq n\Bigr).$
Veronese-type ideals have been extensively studied, as exemplified by [10, 22, 24, 23]. We identify the following notable connection between this class of ideals and the homological shift ideals of powers of a complementary edge ideal.
@@
(Proposition 5.4 and Theorem 5.5) Let $G$ be a tree, and let $I=I_{c}(G)$ be its complementary edge ideal. Then $\operatorname{HS}_{i}(I^{i})$, up to division by a suitable monomial, is a Veronese-type ideal. Furthermore, every Veronese-type ideal can be realized in this manner.
Studying the homological shift algebra of the complementary edge ideal of a cycle is much more difficult. Regarding this topic, our results are as follows:
@@
(Theorem 5.13) Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal. Then, for all integers $1\leq i<n$ and $s\geq 1$, the following equality holds:
$\mathrm{HS}_{i}(I^{s})=\sum_{\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor}\left(\sum_{F\subseteq[n],\ |F|=i-2k}\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}\right).$
Here, we use the convention that $I^{t}=0$ if $t<0$, and ${\bf x}_{F}$ denotes the monomial $\prod_{i\in F}x_{i}$.
By Theorem 1.9, if $s<\left\lfloor\frac{i}{2}\right\rfloor$, then si+k<0s-i+k<0 for all $k\leq\left\lfloor\frac{i}{2}\right\rfloor$, whence $\mathrm{HS}_{i}(I^{s})=0$. This conclusion is consistent with parts (2) and (3) of Theorem 1.3.
Theorem 1.10.
(Corollary 5.14) Let $G$ be a cycle of length $n$ and let $I_{c}(G)$ be its complementary edge ideal, then $\mathrm{HS}_{i}(\mathcal{R}(I_{c}(G)))$ is generated in degree at most iqi-q for all $i>0$. Here, $q:=\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}$.
Theorems 1.7 and 1.10 show that Conjecture 1.1 holds for the complementary edge ideal of a tree or a cycle, while Theorem 0.7 also shows that the conclusion of this conjecture cannot be improved; that is, $\mathrm{HS}_{i}(\mathcal{R}(I))$ cannot be generated in a degree smaller than $i$ in general.
The article is organized as follows: Section 2 provides the definitions and basic facts that will be used throughout the paper. In Section 3, we prove Theorem 1.2 and discuss several of its consequences. In Section 4, we compute the projective dimension of powers of the complementary edge ideals of trees and cycles, and prove Proposition 1.4 and Theorem 1.5. In Section 5, we are devoted to obtaining explicit expressions for the homological shift ideals of powers of the complementary edge ideals of connected graphs, whereby Theorems 1.61.10 are established.
2. Preliminaries and Basic Definitions
This section collects the definitions and basic facts used throughout the paper. We first retain the notation introduced in the Introduction, then introduce additional notation, core concepts, and key lemmas for our study.
2.1. Basic Notation
For a vector $\mathbf{a}=(a_{1},\ldots,a_{n})\in\mathbb{Z}^{n}$, we set $\mathbf{x}^{\mathbf{a}}:=x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$. For a subset $F\subseteq[n]$, we set $\mathbf{x}_{F}:=\prod_{j\in F}x_{j}$ if $F\neq\emptyset$, and $\mathbf{x}_{F}:=1$ if $F=\emptyset$. For positive integers $j\leq k$, let $[j,k]$ denote the set $\{j,j+1,\ldots,k\}$; in particular, $[1,k]$ is abbreviated as $[k]$. We use the symbol $\sqcup$ to denote the disjoint union.
2.2. Core Concepts: Homological Shift Ideals and Algebras
Given the central role of both the homological shift ideal and algebra in this paper, we restate their definitions, drawn respectively from [23] and [15].